Exercise 18-14 (Part Level Submission) Ivanhoe Company manufactures equipment. Ivanhoes products range from simple automated machinery to complex systems containing numerous components. Unit selling prices range from $200,000 to $1,500,000 and are quoted inclusive of installation. The installation process does not involve changes to the features of the equipment and does not require proprietary information about the equipment in order for the installed equipment to perform to specifications. Ivanhoe has the following arrangement with Winkerbean Inc Winkerbean purchases equipment from Ivanhoe for a price of $1,070,000 and contracts with Ivanhoe to install the equipment. Ivanhoe charges the same price for the equipment irrespective of whether it does the installation or not. Using market data, Ivanhoe determines installation service is estimated to have a standalone selling price of $46,200. The cost of the equipment is $599,000 Winkerbean is obligated to pay Ivanhoe the $1,070,000 upon the delivery and installation of the equipment. Ivanhoe delivers the equipment on June 1, 2017, and completes the installation of the equipment on September 30, 2017. The equipment has a useful life of 10 years. Assume that the equipment and the installation are two distinct performance obligations which should be accounted for separately Assuming Ivanhoe does not have market data with which to determine the standalone selling price of the installation services. As a result, an expected cost plus margin approach is used. The cost of installation is $33,200; Ivanhoe prices these services with a 20% margin relative to cost. ▼ (a) How should the transaction price of $1,070,000 be allocated among the service obligations? (Do not round intermediate calculations. Round final answers to 0 decimal places.) Equipment Installation SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT
Answer #1
a
Total fair value of obligation 1116200 =1070000+46200
Equipment 1025712 =1070000*1070000/1116200
Installation 44288 =1070000*46200/1116200